# CodeFuse-CR-Bench: A Comprehensiveness-aware Benchmark for End-to-End Code Review Evaluation in Python Projects

### v1

@@ 1 Introduction
Code Review (CR) is a core practice in modern software development that aims to improve code quality and identify defects through collaborative inspection Bacchelli & Bird (2013). As Large Language Models (LLMs) increasingly automate complex software engineering tasks, their application to CR holds immense promise for improving software quality and developer productivity. However, the development of such sophisticated tools is fundamentally constrained by how we measure their performance. Current benchmarks, while valuable, evaluate models in isolated, decontextualized settings, creating a significant and growing “reality gap” between measured performance and real-world efficacy.
@@ 2.1 CR-related Tasks
Previously, the vast majority of benchmarks and approaches were constructed to improve CR performance Jiang et al. (2025). Tufano et al. Tufano et al. (2021) proposed Trans-Review, which adopted deep learning models to partially automate specific CR tasks. They trained two models to implement two CR sub-tasks: (i) code revision before review and (ii) code revision after review. Based on this prior work, they Tufano et al. (2022) updated the models from deep learning models to the pre-training model T5 and named it T5-Review. The results demonstrated that T5-Review can outperform previous deep learning models for automating CR tasks. Thongtanunam et al. Thongtanunam et al. (2022) proposed AutoTransform, which leverages a Byte-Pair Encoding (BPE) approach to handle new tokens and a Transformer-based Neural Machine Translation architecture to handle long sequences. It can be used in the task of code revision before review. Zhou et al. Zhou et al. (2023) evaluated the above three automatic CR tools and pre-trained models for three processes in CR: code revision before review, review comment generation, and code revision after review. The results show that a general-purpose pre-trained model CodeT5 can outperform other models in most cases. Li et al. Li et al. (2022a) proposed a review comments generator with pre-training models, which is called AUGER. They collected empirical review data from 11 notable Java projects and constructed a dataset of 10,882 code changes to evaluate the performance of the proposed approach. Li et al. Li et al. (2022b) proposed a pre-trained model that utilized four pre-training tasks tailored specifically for the CR scenario, named CodeReviewer. They focused on three key tasks related to CR activities, including code change quality estimation, review comment generation, and code refinement to evaluate the model. They also constructed a high-quality benchmark dataset based on our collected data for these three tasks and conducted comprehensive experiments on it. The experiments demonstrated the SOTA results. Additionally, some LLM-based CR approaches were proposed. Guo et al. Guo et al. (2024) conducted the first empirical study to explore the capabilities of ChatGPT in CR tasks, specifically focusing on automated code revision after reviews. They constructed a new CR dataset with high quality based on the existing benchmark CodeReview Li et al. (2022b). A SOTA CR tool Li et al. (2022b) was selected as a baseline. The research study provided insights into the potential of ChatGPT in automating the CR process.
@@
As an important upstream sub-task of CR, issue-solving has also been extensively studied. The content of issue-solving can provide more complete contexts for CR. Some issue-solving benchmarks had been proposed. Jimenez et al. Jimenez et al. (2024) proposed a benchmark, SWE-Bench, that evaluates LLMs in resolving an issue (typically a bug report or a feature request) submitted to popular Python GitHub repositories. Zan et al. Zan et al. (2025) enlarged the SWE-Bench dataset by adding other programming languages’ issue-solving and named it Multi-SWE-bench. Hu et al. Hu et al. (2024) proposed FAUN-Eval, a benchmark specifically designed to evaluate the fine-grained issue-solving capabilities of LLMs. It can be used to systematically assess LLMs across three distinct tasks: Question-Answer (QA), fault localization, and code editing.
@@ 2.2 LLM for Software Engineering Tasks
Recently, LLMs have demonstrated revolutionary performance improvements in almost all software-engineering-related tasks. Regarding general LLMs, the GPT series Liang et al. (2024), Claude, Gemini, and others had demonstrated powerful code generation, summarization, and program repair through training on large corpora containing code Feng et al. (2024); Cao et al. (2024); Wang et al. (2025b); Zhao et al. (2023); Fan et al. (2023). Specifically, a systematic comparative analysis was conducted on three advanced LLMs, including ChatGPT (O1), DeepSeek (R1), and Gemini (2.0 Flash thinking), for Python code generation, evaluating their performance in correctness, code quality, and computational efficiency. Each of the three LLMs has its own strengths and limitations. Their findings underscored the inherent trade-offs between efficiency, accuracy, and quality in AI-generated code. Sobo et al. Sobo et al. (2025) investigated the effectiveness of LLMs in generating code for Human-Robot Interaction applications. They compared the performance among ChatGPT 3.5, Gemini 1.5 Pro, and Claude 3.5 Sonnet. The study highlighted the rapid advancement in LLM capabilities for specialized programming tasks while also identifying persistent challenges in spatial reasoning and adherence to specific constraints. Zhang et al. Zhang et al. (2024) evaluated the capability of advanced LLMs, including ChatGPT-4 and Claude, in fixing memory corruption vulnerabilities in real-world C/C++ code. They analyzed both the strengths and limitations of LLMs in automated program repair on genuine code. Sun et al. Sun et al. (2025) conducted the examination of prevalent automated evaluation methods for assessing the quality of summaries generated by LLMs and found that the results of the GPT-4 evaluation method are most closely aligned with human evaluation. They also discussed the limitations of LLMs in generating summarization in logic programming languages. All the software-engineering-related tasks mentioned above are highly relevant to CR in technical terms. Therefore, it is possible for LLMs to perform comprehensive CR tasks. In our paper, we select several representative LLMs and evaluate their comprehensive CR capabilities using our constructed benchmark and designed evaluation metrics.
@@ 3.2.1 Repository Selection
To make CodeFuse-CR-Bench more representative, we adopt a strict approach to selecting CR task instances from various open-source repositories. We focus on Python as it is one of the most popular languages on GitHub, possessing a mature and diverse open-source ecosystem. This ensures a rich source of projects with standardized and high-quality CR practices, which is essential for our benchmark’s validity. Specifically, we first search for the top 1,000 starred Python repositories on GitHub [^1] https://github.com/, a Git-based code hosting and collaboration platform. A higher star rating indicates that the repository has better popularity. To ensure the quality of the repository and acquire a large pool of potential CR data, we further sort the projects by the number of PRs and filter out projects with less than 1,500 PRs, referring to Li’s study Li et al. (2022b). Additionally, we only keep the repositories that were maintained in the last year (from 2024-08-15 to 2025-08-15), i.e., the repositories that had commit or PR records over the past year. The process mentioned above aims to keep only active projects and remove repositories that are forked from other repositories, as the PR number is not inherited. This process yielded 230 Python projects that meet our criteria for activity and maturity.
@@ 3.2.2 PR Crawling & Attribute-based Filtering
In this step, we aim to crawl high-quality PR data. To ensure the collected PR data contains high-quality CR, we filter it based on the following attributes: (1) We only include PRs that have at least 1 closing issue reference; (2) We only collect the PR that is merged into the main branch. A “Merged” status indicates that the code changes associated with the PR were accepted and incorporated into its parent repository. We conduct this process by using GitHub GraphQL API [^2] https://docs.github.com/en/graphql. The PRs selected based on these attributes undergo rigorous CR, thereby ensuring that the CR data selected under these PRs is of high quality Zheng et al. (2025).
@@ 3.2.4 Feature Labeling
In this step, we classify the three attributes: problem domain, difficulty, and review effort. Problem domain refers to the category that the PR’s associated issue problem statement belongs to. Referring to SWE-Bench Jimenez et al. (2024), an issue-solving benchmark, there are nine categories of PR problem domains, as summarized in Table 4. Difficulty means how difficult it would be to implement a PR. It can be categorized as low, medium, and high. Review effort means the effort required to review a code change. It is on a scale of 1 to 5. 5 means the most effort. If the repository contains annotations for the above three attributes, we directly utilize the labels from the repository. Otherwise, we employ the LLM-as-a-judge approach Gu et al. (2025). We leverage the Qwen3-235B-A22B model Yang et al. (2025), constructing prompts from the Title, Body, Commit Patch to Review, and Head Commit Message fields in Table 2 as input to classify the three attributes. At last, we get 40,124 CR task instances.
@@ 3.2.5 Manual Selection & Annotation
In this step, we manually select high-quality CR task instances that cover nine types of PR problem domains from the 40,124 CR task instances that have passed the filtering process above to serve as the benchmark. Specifically, we invite two of the authors who have more than 5 years of Python development experience to conduct the selection. Firstly, we find that a non-negligible percentage of review comments, while linked to source code lines in the commit patch, are unlikely to result in code revision in the next round. This kind of review comment is regarded as a noise comment. For example, if a review: (1) that is a simple case (e.g., “looks good to me”, “Thanks!” or “Nice”); (2) only requests formatting changes with no impact on code logic (e.g., “fix indentation”, “add spaces”); (3) requests adding tests (these do not change the code under review); (4) asks for clarification or explanation without suggesting changes (e.g., “please explain”, “what does this do?”); (5) refer previous comments that cannot be identified (e.g., “same as before”, “see above”); (6) only requests adding comments or documentation (e.g., “add Javadoc”, “document this method”); (7) that is difficult to identify (e.g., “At least here it is clear that the equals method of the implementers of TreeNode is important”), it is considered as a noise and irrelevant review comment. Instances that include these comments are not selected in the final benchmark. In addition, we remove instances whose issue problem statement has images, external hyperlinks, references to specific commit SHAs, and references to other pull requests or issues Jimenez et al. (2024). The instance with fewer than 40 words in the problem statement is also removed. They are not considered in our scenario. Based on the above rules, we have filtered out 7,086 instances. After that, referring to SWE-Bench Jimenez et al. (2024) and Multi-SWE-Bench Zan et al. (2025), we design a questionnaire [^3] To be presented. about instance quality assessment and ask the two developers to complete the questionnaire for each instance. This questionnaire covers the following areas: (1) Problem statement and patch alignment; (2) Review scope and comment coverage; (3) Defects identified in the patch; (4) Difficulty and review effort; (5) Overall patch quality and risk; (6) Dataset suitability; and (7) Confidence. Thereby, we select the 601 high-quality instances from 70 projects by manual annotation.
@@ 4.1 Model-base Evaluation
Model-based evaluation is a black-box evaluation designed to leverage the powerful semantic understanding capabilities of LLMs to score and evaluate the quality of CR. It aims to simulate human judgment of CR’s “semantic quality” and “usefulness”. In our paper, we use two kinds of evaluators in model-based evaluation, namely, the reward model and LLM-as-a-judge Son et al. (2024).
@@ 4.1.1 Reward Model
Reward model Yang et al. (2024) is an AI model that learns human preferences. Its core task is: given an input and one or more model outputs (Output/Candidate), it outputs a scalar score (Reward Score) that predicts the degree of human preference for that output (i.e., the extent to which it is perceived as high quality). In order to train a reward model to evaluate comprehensive CR, we need to collect the corresponding training set and design the training strategy.
@@
Training Strategy. Our reward model is built upon the Qwen3-8B Yang et al. (2025) pretrained LLM, where we replace the final LM head layer with a Multi-Layer Perceptron consisting of two linear layers and a ReLU activation function. The first linear layer has dimensions of $(\text{hidden\_size},\text{hidden\_size})$, while the second layer has dimensions of $(\text{hidden\_size},1)$. The final output represents a relevance score indicating how well the CR correlates with the given query and patch.
The training loss comprises two components. The first component follows a similar approach to the reward model training in InstructGPT Ouyang et al. (2022), employing a contrastive learning-style Bayesian Personalized Ranking (BPR) loss Rendle et al. (2012). Specifically, for the same query and patch, we sample positive and negative reviews along with their corresponding code context (reviewed file content as default) and diff hunks to achieve this objective. The loss is formulated as:
@@
Additionally, considering our application scenario of classifying review relevance, we introduce an auxiliary classification objective using binary cross-entropy loss Wang et al. (2025a), that is,
@@
Implementation Details. For efficient training, we employ Low-Rank Adaptation (LoRA) Hu et al. (2022a) as our parameter-efficient fine-tuning method with a rank of 32 and an alpha value of 16, which significantly reduces the number of trainable parameters while maintaining model performance. The model is trained for 1 epoch with a learning rate of 5e55e-5. To accommodate the comprehensive CR context, we set the maximum sequence length to 24,576 tokens, enabling the model to process lengthy code patches and associated reviews effectively.
We utilize Flash Attention 2 Dao (2024) for memory-efficient attention computation, with mixed precision training with bfloat16. The training process incorporates 100 warm-up steps for learning rate scheduling to ensure stable convergence. The distributed training is conducted using DeepSpeed ZeRO Rajbhandari et al. (2020) Stage 2. All experiments are conducted on 32 NVIDIA A100 GPUs with a total effective batch size of 64.
@@ 4.1.2 LLM-as-a-Judge
To evaluate the CR in a general perspective, we also adopt LLM-as-a-judge to evaluate the CR quality. Specifically, we design a prompt that consists of CR information and evaluation criteria to request an LLM to provide a score for the CR quality. Referring to some research about CR usefulness investigation Yang et al. (2023); Turzo & Bosu (2024), we request LLM to provide a score of CR based on four perspectives, that is, Functionality, Quality, Style, and Documentation. For each perspective, we request LLM to analyze five dimensions of Correctness, Relevance, Clarity, Consistency and Language. Each dimension will be provided a score by LLM and then an average overall score will be calculated at last. The detail prompt is shown in Fig. 5. We use OpenAI o3-2025-04-16 OpenAI (2025b) model to achieve LLM-as-a-judge. Thereby, we can obtain a score provided by LLM.

### v2

@@ 1 Introduction
Code Review (CR) is a core practice in modern software development that aims to improve code quality and identify defects through collaborative inspection (Bacchelli & Bird, 2013). As Large Language Models (LLMs) increasingly automate complex software engineering tasks, their application to CR holds immense promise for improving software quality and developer productivity. However, the development of such sophisticated tools is fundamentally constrained by how we measure their performance. Current benchmarks, while valuable, evaluate models in isolated, decontextualized settings, creating a significant and growing “reality gap” between measured performance and real-world efficacy.
@@ 2.1 CR-related Tasks
Previously, the vast majority of benchmarks and approaches were constructed to improve CR performance (Jiang et al., 2025). Tufano et al. (Tufano et al., 2021) proposed Trans-Review, which adopted deep learning models to partially automate specific CR tasks. They trained two models to implement two CR sub-tasks: (i) code revision before review and (ii) code revision after review. Based on this prior work, they (Tufano et al., 2022) updated the models from deep learning models to the pre-training model T5 and named it T5-Review. The results demonstrated that T5-Review can outperform previous deep learning models for automating CR tasks. Thongtanunam et al. (Thongtanunam et al., 2022) proposed AutoTransform, which leverages a Byte-Pair Encoding (BPE) approach to handle new tokens and a Transformer-based Neural Machine Translation architecture to handle long sequences. It can be used in the task of code revision before review. Zhou et al. (Zhou et al., 2023) evaluated the above three automatic CR tools and pre-trained models for three processes in CR: code revision before review, review comment generation, and code revision after review. The results show that a general-purpose pre-trained model CodeT5 can outperform other models in most cases. Li et al. (Li et al., 2022a) proposed a review comments generator with pre-training models, which is called AUGER. They collected empirical review data from 11 notable Java projects and constructed a dataset of 10,882 code changes to evaluate the performance of the proposed approach. Li et al. (Li et al., 2022b) proposed a pre-trained model that utilized four pre-training tasks tailored specifically for the CR scenario, named CodeReviewer. They focused on three key tasks related to CR activities, including code change quality estimation, review comment generation, and code refinement to evaluate the model. They also constructed a high-quality benchmark dataset based on our collected data for these three tasks and conducted comprehensive experiments on it. The experiments demonstrated the SOTA results. Additionally, some LLM-based CR approaches were proposed. Guo et al. (Guo et al., 2024) conducted the first empirical study to explore the capabilities of ChatGPT in CR tasks, specifically focusing on automated code revision after reviews. They constructed a new CR dataset with high quality based on the existing benchmark CodeReview (Li et al., 2022b). A SOTA CR tool (Li et al., 2022b) was selected as a baseline. The research study provided insights into the potential of ChatGPT in automating the CR process.
@@
As an important upstream sub-task of CR, issue-solving has also been extensively studied. The content of issue-solving can provide more complete contexts for CR. Some issue-solving benchmarks had been proposed. Jimenez et al. (Jimenez et al., 2024) proposed a benchmark, SWE-Bench, that evaluates LLMs in resolving an issue (typically a bug report or a feature request) submitted to popular Python GitHub repositories. Zan et al. (Zan et al., 2025) enlarged the SWE-Bench dataset by adding other programming languages’ issue-solving and named it Multi-SWE-bench. Hu et al. (Hu et al., 2024) proposed FAUN-Eval, a benchmark specifically designed to evaluate the fine-grained issue-solving capabilities of LLMs. It can be used to systematically assess LLMs across three distinct tasks: Question-Answer (QA), fault localization, and code editing.
@@ 2.2 LLM for Software Engineering Tasks
Recently, LLMs have demonstrated revolutionary performance improvements in almost all software-engineering-related tasks. Regarding general LLMs, the GPT series (Liang et al., 2024), Claude, Gemini, and others had demonstrated powerful code generation, summarization, and program repair through training on large corpora containing code (Feng et al., 2024; Cao et al., 2024; Wang et al., 2025b; Zhao et al., 2023; Fan et al., 2023). Specifically, a systematic comparative analysis was conducted on three advanced LLMs, including ChatGPT (O1), DeepSeek (R1), and Gemini (2.0 Flash thinking), for Python code generation, evaluating their performance in correctness, code quality, and computational efficiency. Each of the three LLMs has its own strengths and limitations. Their findings underscored the inherent trade-offs between efficiency, accuracy, and quality in AI-generated code. Sobo et al. (Sobo et al., 2025) investigated the effectiveness of LLMs in generating code for Human-Robot Interaction applications. They compared the performance among ChatGPT 3.5, Gemini 1.5 Pro, and Claude 3.5 Sonnet. The study highlighted the rapid advancement in LLM capabilities for specialized programming tasks while also identifying persistent challenges in spatial reasoning and adherence to specific constraints. Zhang et al. (Zhang et al., 2024) evaluated the capability of advanced LLMs, including ChatGPT-4 and Claude, in fixing memory corruption vulnerabilities in real-world C/C++ code. They analyzed both the strengths and limitations of LLMs in automated program repair on genuine code. Sun et al. (Sun et al., 2025) conducted the examination of prevalent automated evaluation methods for assessing the quality of summaries generated by LLMs and found that the results of the GPT-4 evaluation method are most closely aligned with human evaluation. They also discussed the limitations of LLMs in generating summarization in logic programming languages. All the software-engineering-related tasks mentioned above are highly relevant to CR in technical terms. Therefore, it is possible for LLMs to perform comprehensive CR tasks. In our paper, we select several representative LLMs and evaluate their comprehensive CR capabilities using our constructed benchmark and designed evaluation metrics.
@@ 3.2.1 Repository Selection
To make CodeFuse-CR-Bench more representative, we adopt a strict approach to selecting CR task instances from various open-source repositories. We focus on Python as it is one of the most popular languages on GitHub, possessing a mature and diverse open-source ecosystem. This ensures a rich source of projects with standardized and high-quality CR practices, which is essential for our benchmark’s validity. Specifically, we first search for the top 1,000 starred Python repositories on GitHub [^1] https://github.com/, a Git-based code hosting and collaboration platform. A higher star rating indicates that the repository has better popularity. To ensure the quality of the repository and acquire a large pool of potential CR data, we further sort the projects by the number of PRs and filter out projects with less than 1,500 PRs, referring to Li’s study (Li et al., 2022b). Additionally, we only keep the repositories that were maintained in the last year (from 2024-08-15 to 2025-08-15), i.e., the repositories that had commit or PR records over the past year. The process mentioned above aims to keep only active projects and remove repositories that are forked from other repositories, as the PR number is not inherited. This process yielded 230 Python projects that meet our criteria for activity and maturity.
@@ 3.2.2 PR Crawling & Attribute-based Filtering
In this step, we aim to crawl high-quality PR data. To ensure the collected PR data contains high-quality CR, we filter it based on the following attributes: (1) We only include PRs that have at least 1 closing issue reference; (2) We only collect the PR that is merged into the main branch. A “Merged” status indicates that the code changes associated with the PR were accepted and incorporated into its parent repository. We conduct this process by using GitHub GraphQL API [^2] https://docs.github.com/en/graphql. The PRs selected based on these attributes undergo rigorous CR, thereby ensuring that the CR data selected under these PRs is of high quality (Zheng et al., 2025).
@@ 3.2.4 Feature Labeling
In this step, we classify the three attributes: problem domain, difficulty, and review effort. Problem domain refers to the category that the PR’s associated issue problem statement belongs to. Referring to SWE-Bench (Jimenez et al., 2024), an issue-solving benchmark, there are nine categories of PR problem domains, as summarized in Table 4. Difficulty means how difficult it would be to implement a PR. It can be categorized as low, medium, and high. Review effort means the effort required to review a code change. It is on a scale of 1 to 5. 5 means the most effort. If the repository contains annotations for the above three attributes, we directly utilize the labels from the repository. Otherwise, we employ the LLM-as-a-judge approach (Gu et al., 2025). We leverage the Qwen3-235B-A22B model (Yang et al., 2025), constructing prompts from the Title, Body, Commit Patch to Review, and Head Commit Message fields in Table 2 as input to classify the three attributes. At last, we get 40,124 CR task instances.
@@ 3.2.5 Manual Selection & Annotation
In this step, we manually select high-quality CR task instances that cover nine types of PR problem domains from the 40,124 CR task instances that have passed the filtering process above to serve as the benchmark. Specifically, we invite two of the authors who have more than 5 years of Python development experience to conduct the selection. Firstly, we find that a non-negligible percentage of review comments, while linked to source code lines in the commit patch, are unlikely to result in code revision in the next round. This kind of review comment is regarded as a noise comment. For example, if a review: (1) that is a simple case (e.g., “looks good to me”, “Thanks!” or “Nice”); (2) only requests formatting changes with no impact on code logic (e.g., “fix indentation”, “add spaces”); (3) requests adding tests (these do not change the code under review); (4) asks for clarification or explanation without suggesting changes (e.g., “please explain”, “what does this do?”); (5) refer previous comments that cannot be identified (e.g., “same as before”, “see above”); (6) only requests adding comments or documentation (e.g., “add Javadoc”, “document this method”); (7) that is difficult to identify (e.g., “At least here it is clear that the equals method of the implementers of TreeNode is important”), it is considered as a noise and irrelevant review comment. Instances that include these comments are not selected in the final benchmark. In addition, we remove instances whose issue problem statement has images, external hyperlinks, references to specific commit SHAs, and references to other pull requests or issues (Jimenez et al., 2024). The instance with fewer than 40 words in the problem statement is also removed. They are not considered in our scenario. Based on the above rules, we have filtered out 7,086 instances. After that, referring to SWE-Bench (Jimenez et al., 2024) and Multi-SWE-Bench (Zan et al., 2025), we design a questionnaire [^3] To be presented. about instance quality assessment and ask the two developers to complete the questionnaire for each instance. This questionnaire covers the following areas: (1) Problem statement and patch alignment; (2) Review scope and comment coverage; (3) Defects identified in the patch; (4) Difficulty and review effort; (5) Overall patch quality and risk; (6) Dataset suitability; and (7) Confidence. Thereby, we select the 601 high-quality instances from 70 projects by manual annotation.
@@ 4.1 Model-base Evaluation
Model-based evaluation is a black-box evaluation designed to leverage the powerful semantic understanding capabilities of LLMs to score and evaluate the quality of CR. It aims to simulate human judgment of CR’s “semantic quality” and “usefulness”. In our paper, we use two kinds of evaluators in model-based evaluation, namely, the reward model and LLM-as-a-judge (Son et al., 2024).
@@ 4.1.1 Reward Model
Reward model (Yang et al., 2024) is an AI model that learns human preferences. Its core task is: given an input and one or more model outputs (Output/Candidate), it outputs a scalar score (Reward Score) that predicts the degree of human preference for that output (i.e., the extent to which it is perceived as high quality). In order to train a reward model to evaluate comprehensive CR, we need to collect the corresponding training set and design the training strategy.
@@
Training Strategy. Our reward model is built upon the Qwen3-8B (Yang et al., 2025) pretrained LLM, where we replace the final LM head layer with a Multi-Layer Perceptron consisting of two linear layers and a ReLU activation function. The first linear layer has dimensions of $(\text{hidden\_size},\text{hidden\_size})$, while the second layer has dimensions of $(\text{hidden\_size},1)$. The final output represents a relevance score indicating how well the CR correlates with the given query and patch.
The training loss comprises two components. The first component follows a similar approach to the reward model training in InstructGPT (Ouyang et al., 2022), employing a contrastive learning-style Bayesian Personalized Ranking (BPR) loss (Rendle et al., 2012). Specifically, for the same query and patch, we sample positive and negative reviews along with their corresponding code context (reviewed file content as default) and diff hunks to achieve this objective. The loss is formulated as:
@@
Additionally, considering our application scenario of classifying review relevance, we introduce an auxiliary classification objective using binary cross-entropy loss (Wang et al., 2025a), that is,
@@
Implementation Details. For efficient training, we employ Low-Rank Adaptation (LoRA) (Hu et al., 2022a) as our parameter-efficient fine-tuning method with a rank of 32 and an alpha value of 16, which significantly reduces the number of trainable parameters while maintaining model performance. The model is trained for 1 epoch with a learning rate of 5e55e-5. To accommodate the comprehensive CR context, we set the maximum sequence length to 24,576 tokens, enabling the model to process lengthy code patches and associated reviews effectively.
We utilize Flash Attention 2 (Dao, 2024) for memory-efficient attention computation, with mixed precision training with bfloat16. The training process incorporates 100 warm-up steps for learning rate scheduling to ensure stable convergence. The distributed training is conducted using DeepSpeed ZeRO (Rajbhandari et al., 2020) Stage 2. All experiments are conducted on 32 NVIDIA A100 GPUs with a total effective batch size of 64.
@@ 4.1.2 LLM-as-a-Judge
To evaluate the CR in a general perspective, we also adopt LLM-as-a-judge to evaluate the CR quality. Specifically, we design a prompt that consists of CR information and evaluation criteria to request an LLM to provide a score for the CR quality. Referring to some research about CR usefulness investigation (Yang et al., 2023; Turzo & Bosu, 2024), we request LLM to provide a score of CR based on four perspectives, that is, Functionality, Quality, Style, and Documentation. For each perspective, we request LLM to analyze five dimensions of Correctness, Relevance, Clarity, Consistency and Language. Each dimension will be provided a score by LLM and then an average overall score will be calculated at last. The detail prompt is shown in Fig. 5. We use OpenAI o3-2025-04-16 (OpenAI, 2025b) model to achieve LLM-as-a-judge. Thereby, we can obtain a score provided by LLM.